• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Armstrong

    Log entry time 20:39:53 on September 13, 2009

    Entry number 289401

    This entry is a followup to: 289366

    keyword=low beam cut readjusted for 100 muA

    We are back at 100 muA and the beam looks stable. I have re-edited the low-beam threshold cut for the PanGUIN to be 20000 (was 16000 for 80 muA running). The file I edited was ~/db/cuts/happex3_oversample1.cuts (not (~/db/cut/happex3_oversample2.cuts) from Kent's entry, I think that might have been a typo. But maybe I am on the wrong account/machine (I am on apar/adaql6)